OpenStudy (anonymous):

Be f(x)=e^3x-1 evaluate for f(Ln2)

OpenStudy (anonymous):

replace the x in the equation by ln2

OpenStudy (anonymous):

now you have \[f(\ln2)=e ^{3\ln2}-1\]

OpenStudy (anonymous):

I did it. Thats all?

OpenStudy (anonymous):

Tyvm Esamalla but why -1 is now on the bottom

OpenStudy (anonymous):

is your equation \[e ^{3x}-1 or e ^{3x-1}\]

OpenStudy (anonymous):

\[f(\ln2)=e ^{3\ln2-1}\]

OpenStudy (anonymous):

this is the ans if your equation is the second

OpenStudy (anonymous):

\[f(x)=e ^{3x-1}=e ^{3x}e ^{-1}=\frac{ e ^{3x} }{ e }\] \[f \left( \ln 2 \right)=\frac{ e ^{3\ln 2} }{e }=\frac{e ^{\ln 2^{3}} }{ e }=\frac{ 8 }{e }\]